OpenStudy (anonymous):

Solve: 7(x − 1) = 5(x + 2) + 1 9 18 2 4

OpenStudy (anonymous):

this question will have to answers. 5x-7/x-1=x+1 ( x+1 )( x-1 ) = 5x-7 x^2 - 1 = 5x - 7 x^2-5x+6= 0 Delta = b^2 - 4ac= 25 - 4 ( 1^2 )( 6 )= 25 - 24 = 1 x = ( 5 + 1 )/2= 3 x = ( 5 -1 )/2 = 2